OpenStudy (anonymous):

2m^2+km-3 find all intergers of k so that the trinomial can be factored.

OpenStudy (anonymous):

2*-3=-6 3-2=1 k=1,-1 6-1=5 6*-1=-6 k=5,-5

OpenStudy (anonymous):

thank you so much! im ust confused as to how you did that? could you possibly explain?

OpenStudy (anonymous):

2*-3=-6 k=a-b so that a*b=-6

OpenStudy (anonymous):

but k is b?

OpenStudy (anonymous):

like how would i solve 7m^2+km-5 because i still dont understand it at all...

OpenStudy (anonymous):

7*-5=35 thenk=34 or -34

OpenStudy (anonymous):

correction above 7*-5=-35 a=35,b=-1,35*-1=-35 a=-35,b=1,-35*1=-35

OpenStudy (anonymous):

if the product of a*c=-ve,then always subtract if the product of a*c=+ve,then always break b

OpenStudy (anonymous):

okay that makes a ittle more sense. thank you

OpenStudy (anonymous):

i am talking of ax^2+bx+c=0

OpenStudy (anonymous):

in the above example k=2,-2 or k=34,-34 7-5=2,7*-5=-35 -7+5=-2,-7*5=-35 35-1=34,35*-1=-35 -35+1=-34,-35*1=-35

OpenStudy (anonymous):

where did the 1 come from when you got 34?

OpenStudy (anonymous):

make factors of35 

OpenStudy (anonymous):

35-1=34 -35+1=-34 7-5=2 -7+5=-2